You are going to value Lauryns Doll Co. using the FCF model. After consulting various sources, you find that Lauryns has a
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

We know the fomula for FCF as

FCF = EBIT*(1-tax)+Depreciation-capex - NWC

=> FCF = $57*(1-0.40)+$5.7-$4.3 - 2.6

FCF = $33 million.

Equity beta Asset beta = {1+((1-Tax rate)x)

=> Beta = 1.7/(1+(1-0.4)*0.5) = 1.307

=> Discount Rate = Risk Free Rate + (Beta x Market Risk Premium) = 6 + 1.307 * 12 = 21.69%

Now, Firm Value = FCF (1+g) / (R-G)

=> Firm Value = 33 (1+0.02) / (0.2169 - 0.02)

Firm Value = $170.95 million